Circular No. 9088

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 U5 (GRAUER) A. D. Grauer reports the discovery of a comet on four co-added 60-s

CCD exposures taken with the Mount Lemmon 1.5-m reflector, the object having a very condensed coma ∼ 7′′ wide and a faint tail 20′′ long in p.a. 280^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, W. H. Ryan reported finding a tail in p.a. ∼ 270^{o} on CCD exposures taken by E. V. Ryan and himself with the Magdalena Ridge Observatory 2.4-m f/8.9 reflector on Oct. 24.5 and 25.4 UT.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 23.45917 8^{h}57^{m}56.^{s}19 +18^{o}02′13′′.0 19.4 Grauer

Additional astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-U100:

T = 2010 Aug. 23.034 TT  ω = 139.^{o}324
Ω = 118.302 }2000.0
q = 0.57126 AU  i = 43.016

COMETS C/2009 J3–J10 (SOHO) Further to IAUC 9084, additional Kreutz sungrazers have been found

on SOHO website images. C/2009 J3 (peak mag ∼ 6.5), C/2009 J5 (mag ∼ 8–8.5), C/2009 J7 (mag ∼ 8.5), and C/2009 J9 (mag ∼ 7.5–8) were all stellar in appearance and generally small or tiny. C/2009 J4 was a diffuse streak (mag ∼ 7.5–8). C/2009 J6 was slightly diffuse and tiny (mag ∼ 7.5) in C3 images, and slightly diffuse but elongated in C2 images. C/2009 J8 was quite condensed (mag ∼ 7). C/2009 J10 was quite diffuse and slightly elongated (mag ∼ 7.5–8).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 J3 May 3.846 2^{h}51.^{m}6 +13^{o}05′ C3  TH 2009-N46
C/2009 J4  5.033 2 50.3 +14 41 C2  MU 2009-N46
C/2009 J5  6.192 2 56.8 +15 05 C2  BZ 2009-N46
C/2009 J6  6.529 3 00.7 +14 54 C3/2 RK 2009-N47
C/2009 J7  6.617 2 58.8 +15 15 C2  ML 2009-N47
C/2009 J8  8.633 3 06.2 +15 50 C2  MK 2009-N47
C/2009 J9  9.576 3 09.7 +15 59 C2  MK 2009-N47
C/2009 J10  11.351 3 16.1 +16 26 C2  RK 2009-N49
2009 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green